if the unbalance force on the ball is 7.25*10^-2n and 4.35*10^-2j of work is done to lift the ball how far does the catcher lift the ball?
1
Expert's answer
2012-02-09T09:54:39-0500
by defenition of work A=F*d, where A - is work, F - forse, d - distance. Then we
have: d=A/F=4.35*10^-2j/7.25*10^-2n=0,6 metres.